# DAISI: Data Assimilation with Inverse Sampling using Stochastic Interpolants

### v1

@@ 3.2 Complexity Analysis
Assuming the dynamics model $\mathcal{F}(\cdot)$ incurs a cost of $\mathcal{O}(f(d_{x}))$, the prediction step (12) costs $\mathcal{O}(Jf(d_{x}))$. Denoting by $\mathcal{O}(g(d_{x},d_{y}))$ the cost of a single Euler–Maruyama (EM) step for solving (6) or (7) (with or without guidance), the analysis step costs $\mathcal{O}(JTg(d_{x},d_{y}))$, where $T$ is the number of EM steps. Hence, the overall computational complexity of DAISI is $\mathcal{O}\big(NJ(f(d_{x})+Tg(d_{x},d_{y}))\big)$. For comparison, the ensemble transform Kalman filter (ETKF) has cost $\mathcal{O}\big(NJ(f(d_{x})+(d_{x}+d_{y})J+J^{2})\big)$. Since the forward evaluation of the U-net parameterizing the drift ${\bm{b}}(t,{\bm{z}})$ scales linearly with input dimension (i.e., $g(d_{x},d_{y})=\mathcal{O}(d_{x}+d_{y})$), DAISI achieves comparable cost to ETKF when $T\sim J$, while scaling linearly in ensemble size JJ—in contrast to ETKF’s cubic scaling in $J$. However, in settings where $J$ is small (e.g., $\mathcal{O}(10)$), DAISI can be more expensive due to the need for $T=\mathcal{O}(10^{2})$ EM steps to accurately solve the forward and backward SDEs.
@@ 3.3 Related Works
Early work on diffusion priors for data assimilation includes Score-based Data Assimilation (SDA) (rozet2023score), which learns the score of full trajectories to perform smoothing without an explicit forecast model. While effective, it is memory-intensive and scales poorly for filtering.
Closer to our setting, Yang et al. (yang2025generative) address linear filtering by guiding a pre-trained diffusion model using both forecasts and observations. Instead of solving the backward SDE, they rely on SDEdit (meng2022sdedit), which partially inverts a sample by re-noising it, and incorporate observations via RePaint (lugmayr2022repaint), an iterative noising-denoising procedure. Together, these steps create a trade-off between preserving forecast information and enforcing the conditioning. In Appendix D.4.6, we show that replacing the backward SDE with SDEdit leads to substantially worse results.
The Ensemble Score Filter (EnSF) (bao2024ensemble) replaces the learned score with an analytical approximation, enabling training-free assimilation of nonlinear observations. However, it struggles in sparse settings due to the lack of a learned structure. An extension of EnSF in latent-space (si2025latentensf) partially addresses this; however, this does not easily amortize over observation models and its results depend heavily on the quality of the trained autoencoder. EnSF and its variants have been applied to the SQG model (bao2025nonlinear; yin2024scalable; liang2025ensemble), and extensions with Gaussian mixture score approximations (zhang2025iensf) and problem-dependent data coupling (transue2025flow) have been proposed.
Finally, recent works explore guidance for diffusion-based forecasting models. For instance, FlowDAS (chen2025flowdas) employs stochastic interpolants, while (savary2025training) uses guidance and particle filters to steer the GenCast model (gencast). These methods depend heavily on the accuracy of the learned one-step forecast distribution $p({\bm{x}}_{n+1}|{\bm{x}}_{n})$. When this distribution is too concentrated, the model effectively “expects” the next state to lie within a narrow range, and struggles to correct toward observations that fall outside it. Because DAISI learns the marginal distribution instead, it is less prone to such failures.
@@ 4 Experiments
We evaluate DAISI on three systems: the Lorenz ’63 (L63) system (lorenz1963deterministic), a Surface Quasi-Geostrophic system (SQG) (tulloch2009quasigeostrophic), and SEVIR, a real-world radar dataset (veillette2020sevir).
@@ 4.1 Lorenz ’63 System
We evaluate DAISI on the Lorenz ’63 (L63) system to assess its filtering performance. As discussed in Section 3.1, DAISI does not sample exactly from the filtering distribution; our goal here is to demonstrate how an appropriate choice of $t_{\min}$ and $\epsilon$ can substantially reduce errors. The L63 system is integrated with standard parameters for 5000 RK4 time steps from random initial conditions, and the final 500 steps are used to test the assimilation methods. For reference, we compare DAISI with the Bootstrap Particle Filter (BPF) (gordon1993novel), a gold standard for low-dimensional filtering. As BPF is asymptotically exact, we take this as the “ground truth” filtering result to compare against. We also use an asymptotically exact guidance method for DAISI (see Appendix A.5.3) to isolate DAISI’s intrinsic error. For further details on the experimental setup, see Appendix D.3.
Table 1 summarizes the RMSE and probabilistic metrics—the Continuous Ranked Probability Score (CRPS) and Spread–Skill Ratio (SSR). For comparison, we also include the results of DAISI without inverse-sampling; this performs substantially worse than BPF across all metrics, demonstrating that incorporating dynamical information is essential for accurate filtering. With the default hyperparameters $t_{\min}\approx 0$ and $\epsilon=0$, DAISI still underperforms, and yields worse results than the version without inversion. Tuning $t_{\min}$ alone [^1] The hyperparameters are tuned on a short held-out validation trajectory to minimize CRPS. already leads to substantial improvement, and jointly tuning $t_{\min}$ and $\epsilon$ produces the best performance, achieving metrics closest to those of BPF.
@@ 4.2 Surface Quasi-Geostrophic (SQG) Dynamics
We next evaluate DAISI on a two-layer Surface Quasi-Geostrophic (SQG) model, a standard benchmark for turbulent geophysical flows. The system evolves potential temperature on two horizontal surfaces confining a fluid. Despite its simplicity, SQG exhibits strong sensitivity to initial conditions and exhibit multiscale turbulent behavior while remaining computationally tractable, making it a valuable benchmark for DA (tulloch2009quasigeostrophic).
Following (liang2025ensemble), we use a $64\times 64$ grid over 100 time steps with 3-hour intervals, and additionaly include a single $256\times 256$ experiment to assess scalability. DAISI is tested across a range of observation operators, noise levels, and sparsity settings listed in Table 2. Since the guidance methods in Section 2.3 have been applied almost exclusively to Gaussian observations, we stay within this setting, but note that this is not a limitation with DAISI itself.
We compare DAISI to both classical and ML-based DA methods. Classical baselines include the Localized Ensemble Transport Filter (LETKF), while ML baselines include FlowDAS and the Ensemble Score Filter (EnSF). DAISI uses a U-Net architecture based on Karras2022edm with 3.5M parameters. With the exception of FlowDAS, we initialize all methods from ${\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0}^{\mathrm{gt}},\sigma_{\mathrm{init}}^{2}\mathbf{I})$, $\sigma_{\mathrm{init}}=3$, and propagate ensemble members using the numerical model. FlowDAS instead uses its own learned autoregressive forecast model conditioned on the six previous states. Therefore assimilation begins from samples drawn from $p({\bm{x}}_{0}\mid{\bm{x}}_{-1}^{\mathrm{gt}},\dots,{\bm{x}}_{-6}^{\mathrm{gt}})$. All methods use 20 members.
@@
SEVIR $\mathbf{0.016}{\scriptstyle\pm 0.01}$ - $\underline{0.045}{\scriptstyle\pm 0.01}$ -
Table 3 summarizes CRPS across all SQG experiments. DAISI consistently achieves accurate assimilation, producing temporally coherent and physically plausible ensemble members (Figures 4 and 5). It matches LETKF in the noisy and high-dimensional settings, and clearly outperforms it under sparse or nonlinear observations. DAISI also remains stable when assimilation is performed every 12 hours instead of every 3, whereas LETKF degrades (Figure 12). In multimodal settings, DAISI reliably tracks all plausible modes, whereas LETKF collapses to a single mode and typically diverges once that mode becomes inconsistent (Figure 21).
For the saturating arctan observations, DAISI performs comparably to EnSF, which handles such nonlinearities well. However, EnSF fails in other regimes due to the lack of a learned prior, and is prone to mode collapse, requiring inflation to perform well. LETKF diverges under these nonlinear observations, and FlowDAS underperforms in all settings, reflecting its sensitivity to observations outside the support of its forecast model. Replacing the numerical model with the FlowDAS forecast model within DAISI shows no performance degradation, confirming FlowDAS’s failures arise from the filtering scheme rather than its forecast model (Table 11). Additional results are provided in Appendix D.4.5.
@@ 4.3 Precipitation Nowcasting using SEVIR
To evaluate on a real-world dataset, we apply DAISI to the Storm EVent Imagery and Radar (SEVIR) dataset (veillette2020sevir), a radar observation dataset of convective storms over the United States. We use the vertically integrated liquid on a $384\times$$384\text{\,}\mathrm{km}$ grid at $2\text{\,}\mathrm{km}$ resolution available every $10\text{\,}\mathrm{min}$ for $250\text{\,}\mathrm{min}$ (gao2023prediff).
@@
This dataset has been applied to DA using the FlowDAS method in (chen2025flowdas). Thus, we mirror their setting and consider a linear Gaussian observation with standard deviation 0.001 and 10% sparsity. To generate the forecasts, we use the pre-trained forecasting model of (chen2025flowdas). This takes six consecutive frames as input and predicts the state 10 min into the future.
@@
As shown in Figure 5, both DAISI and FlowDAS are able to accurately reconstruct the state, although the extreme events are better represented in DAISI. This is also reflected by the much lower CRPS, as shown in Table 3.
5 Conclusion
@@
We introduced DAISI, a robust and flexible filtering framework built on flow-based generative priors. DAISI avoids retraining at every assimilation step by leveraging a fixed, pre-trained flow model and enables zero-shot conditional sampling for arbitrary observation operators. A key component is the inversion of the generative SDE: running the flow backward from the forecast ensemble recovers latent states that serve as informative initialization for conditional sampling, effectively combining the prior, the forecast model, and observational information in a unified way.
Empirically, DAISI delivers accurate filtering results across a spectrum of challenging settings, including sparse, noisy, nonlinear, and multimodal observations. It maintains physical coherence over long horizons, remains stable under infrequent assimilation, and outperforms classical approaches where they break down. Its modular design further supports seamless integration with both numerical simulators and modern ML-based forecasting models.
Overall, DAISI offers a flexible, robust, and expressive probabilistic data assimilation framework, capable of capturing complex posterior structures in high dimensions that are beyond the reach of classical filters.
6 Limitations & Future Work
A fundamental limitation of DAISI is that it does not sample exactly from the true filtering distribution. While our experiments show that tuning $t_{\min}$ and $\epsilon$ can mitigate this discrepancy, developing principled correction schemes for debiasing remains an important direction for future work.
DAISI also inherits the limitations of the underlying generative model. For the $64\times 64$ SQG experiment, our trained generative prior underestimated power across scales, causing degraded performance for the DAISI variant without inversion. Notably, DAISI itself appears robust to this issue, suggesting that the inversion step helps counteract spectral biases of the generative model; understanding this corrective effect more formally is an interesting open question.
Another limitation concerns the guidance mechanism. In this work, we use gradient-based guidance, which is known to introduce bias. Certain iterative schemes that repeatedly inject noise into the dynamics (e.g., (zhang2025improving)) are unlikely to work well in combination with DAISI.
@@ 6 Limitations & Future Work
From a computational standpoint, DAISI inherits the high sampling cost of flow-based generative models, driven by the high number of function evaluations required to integrate generative SDEs. Performing DA in latent spaces (andry2025appa) may offer avenues for reducing these costs.
Finally, the role of $\epsilon$ is richer than the limited setting considered here. More expressive parameterizations—such as matrix-valued or ensemble-adaptive noise levels—may offer additional flexibility and improved performance. We leave these extensions for future investigation.
@@ A.1 Stochastic Interpolants
where ${\bm{z}}_{1}\sim\rho_{1}$ and ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. We have the following result, found in Theorem 2.6 of (albergo2023stochastic).
@@
Furthermore, (albergo2023stochastic) shows that the drift ${\bm{b}}(t,{\bm{z}})$ can be learned by minimizing the objective

### v2

@@ 3.2 Complexity Analysis
Assuming the dynamics model $\mathcal{F}(\cdot)$ incurs a cost of $\mathcal{O}(f(d_{x}))$, the prediction step (12) costs $\mathcal{O}(Jf(d_{x}))$. Denoting by $\mathcal{O}(g(d_{x},d_{y}))$ the cost of a single Euler–Maruyama (EM) step for solving (6) or (7) (with or without guidance), the analysis step costs $\mathcal{O}(JTg(d_{x},d_{y}))$, where $T$ is the number of EM steps. Hence, the overall computational complexity of DAISI is $\mathcal{O}\big(J(f(d_{x})+Tg(d_{x},d_{y}))\big)$. For comparison, the ensemble transform Kalman filter (ETKF) has cost $\mathcal{O}\big(J(f(d_{x})+(d_{x}+d_{y})J+J^{2})\big)$. Since the forward evaluation of the U-net parameterizing the drift ${\bm{b}}(t,{\bm{z}})$ scales linearly with input dimension (i.e., $g(d_{x},d_{y})=\mathcal{O}(d_{x}+d_{y})$), DAISI achieves comparable cost to ETKF when $T\sim J$, while scaling linearly in ensemble size JJ—in contrast to ETKF’s cubic scaling in $J$. However, in settings where $J$ is small (e.g., $\mathcal{O}(10)$), DAISI can be more expensive due to the need for $T=\mathcal{O}(10^{2})$ EM steps to accurately solve the forward and backward SDEs.
@@ 3.3 Related Works
Early work on diffusion priors for data assimilation includes Score-based Data Assimilation (SDA) (Rozet and Louppe, 2023), which learns the score of full trajectories to perform smoothing without an explicit forecast model. While effective, it is memory-intensive and scales poorly to the size of the filtering window.
Closer to our setting, Yang et al. (2025) address linear filtering by guiding a pre-trained diffusion model using both forecasts and observations. Instead of solving the backward SDE, they rely on SDEdit (Meng et al., 2022), which partially inverts a sample by re-noising it, and incorporate observations via RePaint (Lugmayr et al., 2022), an iterative noising-denoising procedure. Together, these steps create a trade-off between preserving forecast information and enforcing the conditioning. In Appendix D.4.6, we show that replacing the backward SDE with SDEdit leads to substantially worse results.
The Ensemble Score Filter (EnSF) (Bao et al., 2024b) replaces the learned score with an analytical approximation, enabling training-free assimilation of nonlinear observations. However, it struggles in sparse settings due to the lack of a learned structure. An extension of EnSF in latent-space (Si and Chen, 2025) partially addresses this; however, this does not easily amortize over observation models and its results depend heavily on the quality of the trained autoencoder. EnSF and its variants have been applied to the Surface Quasi-Geostrophic model (Bao et al., 2025; Yin et al., 2024; Liang et al., 2025), and extensions with Gaussian mixture score approximations (Zhang et al., 2025b) and problem-dependent data coupling (Transue et al., 2025) have been proposed.
Finally, recent works explore guidance for diffusion-based forecasting models. For instance, FlowDAS (Chen et al., 2025) employs stochastic interpolants to learn a one-step forecast distribution $p({\bm{x}}_{n+1}|{\bm{x}}_{n})$, which can serve as a generative prior to condition on new observations ${\bm{y}}_{n+1}$ via guidance. However, this procedure effectively targets the local conditional $p({\bm{x}}_{n+1}|{\bm{x}}_{n},{\bm{y}}_{n+1})$, which can be arbitrarily far from the true filtering distribution $p({\bm{x}}_{n+1}|{\bm{y}}_{1:n+1})$. The work (Savary et al., 2025) [^1] A minor difference from (Chen et al., 2025) is that it uses GenCast (Price et al., 2025) as the predictive model instead of the stochastic interpolant. proposes a correction to this mismatch using particle reweighting. However, this reintroduces particle-filter-style weight degeneracy and thus susceptibility to the curse of dimensionality.
@@ 4 Experiments
We evaluate DAISI on three systems: the Lorenz ’63 (L63) system (Lorenz, 1963), a Surface Quasi-Geostrophic system (SQG) (Tulloch and Smith, 2009), and SEVIR, a real-world radar dataset (Veillette et al., 2020).
@@ 4.1 Lorenz ’63 System
The goal of this experiment is to demonstrate on the L63 system that an appropriate choice of $t_{\min}$ and $\epsilon$ can approximate the filtering distribution closely. The L63 system is integrated with standard parameters from random initial conditions, and the final 500 steps are used to test the assimilation methods. For reference, we compare DAISI with the Bootstrap Particle Filter (BPF) (Gordon et al., 1993). As BPF is asymptotically exact and non-degenerate for the low-dimensional L63 system, we take this as the “ground truth” filtering result to compare against. We also use an asymptotically exact guidance method for DAISI (see Appendix A.5.3) to isolate DAISI’s intrinsic error. For further details on the experimental setup, see Appendix D.3.
Table 1 summarizes the RMSE and probabilistic metrics—the Continuous Ranked Probability Score (CRPS) and Spread–Skill Ratio (SSR). For comparison, we also include the results of DAISI without inverse-sampling (i.e., sampling from $\mathbb{P}_{\infty}^{\bm{y}}$); this performs substantially worse than BPF across all metrics, demonstrating that incorporating dynamical information is essential for accurate filtering. With $t_{\min}\approx 0$ and $\epsilon=0$, DAISI still underperforms, and yields worse results than the version without inversion. Tuning $t_{\min}$ alone [^2] The hyperparameters are tuned on a short held-out validation trajectory to minimize CRPS. already leads to substantial improvement, and jointly tuning $t_{\min}$ and $\epsilon$ produces the best performance, achieving metrics closest to those of BPF.
@@ 4.2 Surface Quasi-Geostrophic (SQG) Dynamics
We next evaluate DAISI on a two-layer Surface Quasi-Geostrophic (SQG) model, a standard benchmark for turbulent geophysical flows. The system evolves potential temperature on two horizontal surfaces confining a fluid. Despite its simplicity, SQG exhibits strong sensitivity to initial conditions and exhibit multiscale turbulent behavior while remaining computationally tractable, making it a valuable benchmark for DA (Tulloch and Smith, 2009).
Following (Liang et al., 2025), we use a $64\times 64$ grid over 100 time steps with 3-hour intervals, and additionally include a single $256\times 256$ experiment to assess scalability. DAISI is tested across a range of observation operators, noise levels, and sparsity settings listed in Table 2. Since the guidance methods in Section 2.3 have been applied almost exclusively to Gaussian observations, we stay within this setting, but note that this is not a limitation with DAISI.
We compare DAISI to both classical and ML-based DA methods. Classical baselines include the Local Ensemble Transform Kalman Filter (LETKF), while ML baselines include FlowDAS and the Ensemble Score Filter (EnSF). DAISI uses a U-Net architecture based on Karras et al. (2022) with 3.5M parameters. With the exception of FlowDAS, we initialize all methods from ${\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0}^{\mathrm{gt}},\sigma_{\mathrm{init}}^{2}\mathbf{I})$, $\sigma_{\mathrm{init}}=3$, and propagate ensemble members using the numerical model. FlowDAS instead uses its own learned autoregressive forecast model conditioned on the six previous states. Therefore assimilation begins from samples drawn from $p({\bm{x}}_{0}\mid{\bm{x}}_{-1}^{\mathrm{gt}},\dots,{\bm{x}}_{-6}^{\mathrm{gt}})$. All methods use 20 members and are tuned specifically for each setting.
@@
SEVIR $\textbf{0.016}{\scriptstyle\pm 0.01}$ $\underline{0.018}{\scriptstyle\pm 0.01}$ $045{\scriptstyle\pm 0.01} 0.0$$075{\scriptstyle\pm 0.02}$
Table 3 summarizes CRPS across all SQG experiments. DAISI consistently achieves accurate assimilation, producing temporally coherent and physically plausible ensemble members (Figures 4 and 5). It matches LETKF in the noisy setting, and clearly outperforms it under sparse or nonlinear observations. DAISI also remains stable when assimilation is performed every 12 hours instead of every 3, whereas LETKF degrades (Figure 12). In multimodal settings, DAISI reliably tracks many plausible modes, whereas LETKF collapses to a single mode and typically diverges if that mode becomes inconsistent. In the high-dimensional experiment, both DAISI and LETKF track the ensemble mean accurately, but exhibit qualitatively different behaviors. DAISI produces smoother reconstructions, while LETKF tends to introduce spurious fine-scale structure. This behavior is likely due to the challenging observation setup and the lack of tuning at this resolution.
For the saturating arctan observations, DAISI performs comparably to EnSF, which handles such nonlinearities well. However, EnSF fails in other regimes due to the lack of a learned prior, and is prone to mode collapse, requiring inflation to perform well. LETKF diverges under these nonlinear observations, and FlowDAS underperforms in all settings, as it samples from $p({\bm{x}}_{n+1}|{\bm{x}}_{n},{\bm{y}}_{n+1})$ rather than $p({\bm{x}}_{n+1}|{\bm{y}}_{1:n+1})$ at each step, leading to errors accumulating over time. Replacing the numerical model with the FlowDAS forecast model within DAISI shows no performance degradation, confirming FlowDAS’s failures arise from the filtering scheme rather than its forecast model (Table 14). Additional results, including RMSE and SSR scores, are provided in Appendix D.4.5. We also display results for more realistic observations simulating moving satellite tracks (refered to as Non-stationary), showing comparable performance to LETKF.
@@ 4.3 Precipitation Nowcasting using SEVIR
To evaluate on a real-world dataset, we apply DAISI to the Storm EVent Imagery and Radar (SEVIR) dataset (Veillette et al., 2020), a radar observation dataset of convective storms over the United States. We use the vertically integrated liquid on a $384\times$$384\text{\,}\mathrm{km}$ grid at $2\text{\,}\mathrm{km}$ resolution available every $10\text{\,}\mathrm{min}$ for $250\text{\,}\mathrm{min}$ (Gao et al., 2023).
@@
This dataset has been applied to DA using the FlowDAS method by Chen et al. (2025). Thus, we mirror their setting and consider a linear Gaussian observation with standard deviation 0.001 and 10% sparsity. To generate the forecasts, we use the pre-trained forecasting model of (Chen et al., 2025). This takes six consecutive frames as input and predicts the state 10 min into the future.
@@
As shown in Figure 5, both DAISI and FlowDAS are able to accurately reconstruct the state, although the peaks are better represented in DAISI. This is also reflected by the much lower CRPS, as shown in Table 3.
4.4 Comparison with Score-based Data Assimilation
We also compare against score-based data assimilation (SDA) (Rozet and Louppe, 2023), which is another generative data assimilation method that uses a pre-trained spatiotemporal diffusion model to condition on all observations—from the past to the present—simultaneously. This makes SDA essentially a smoothing algorithm and therefore not a fully like-for-like baseline for our filtering setting. Hence, we also propose a filtering variant that uses a sliding window for iterative conditioning, which we describe in Appendix D.4.3. We compare against both the filtering and smoothing variants, and display their RMSE, CRPS and SSR scores in Tables 1113. On the linear problems (Noisy, Sparse and SEVIR), SDA smoothing consistently outperforms DAISI. However, this is likely due to the effect of smoothing, which improves past state estimation. Evidently, the results at the last time step are similar to DAISI (Figures 21, 22 and 26). The filtering variant of SDA is slightly worse than DAISI, except on Noisy. On the nonlinear problems (Multimodal and Saturating), both the smoothing and filtering variants of SDA struggle to achieve strong performance.
5 Conclusion and Outlook
@@
We introduced DAISI, a robust and flexible filtering framework built on flow-based generative priors. A key component is the inversion of the generative SDE: running the flow backward from the forecast ensemble recovers latent representations that serve as informative initialization for conditional sampling, effectively combining the prior, the forecast model, and observational information in a unified way. Empirically, DAISI delivers accurate filtering results across a spectrum of challenging settings, including sparse, noisy, nonlinear, and multimodal observations.
@@
Current limitations of DAISI are that firstly, it relies on gradient-based guidance methods during inference, which may be less compatible with iterative schemes that repeatedly inject noise into the dynamics (e.g., (Zhang et al., 2025a)). Secondly, DAISI inherits the high inference cost of ODE/SDE-based generative models, driven by the high number of function evaluations required to integrate generative SDEs. Performing DA in latent spaces (Andry et al., 2025) may offer avenues for reducing these costs. We leave these extensions for future investigation.
@@ A.1 Stochastic Interpolants
where ${\bm{z}}_{1}\sim\rho_{1}$ and ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. We have the following result, found in Theorem 2.6 of (Albergo et al., 2023).
@@
Furthermore, (Albergo et al., 2023) shows that the drift ${\bm{b}}(t,{\bm{z}})$ can be learned by minimizing the objective